John wants to buy a ring with sapphire( Al2O3 in its purest form) that contains 2.22 x 10^22 molecules of Oxygen. What volume of Al2O3 would he need in cm^3? The molar mass of Al2O3 is 101.96 g/mol and the density is 3.95 g/cm^3.
1
Expert's answer
2020-09-14T08:27:14-0400

From the number of Oxygen particles the number of moles may be found:

n=N/Na, where Na is Avogadro's constant.

n=2.22×1022/6.02×1023=0.0369 (moles).

There are 3 atoms of Oxygen in each molecule of Al2O3, so it's number of moles is 3 times lower:

n(Al2O3)=0.0369/3=0.0123(moles).

From molar mass mass of sapphire may be found:

m=n×M.

m=0.0123×101.96=1.2541(g).

From density the needed volume may be found:

V=m/d.

V=1.2541/3.95=0.3175(cm3).

Answer: 0.3175 cm3.